# A Survey of Privacy-Preserving Model Explanations: Privacy Risks, Attacks, and Countermeasures

### v1

@@ 5.2. Defences with Privacy-Preserving SHAP
Several studies have focused on preserve the privacy of users from explanation using Shapley values, including quatization, dimension reduction, multi-party computation, federated learning (Liu et al., 2023b; Liu et al., 2024a), and differential privacy (see Fig. 15).
@@
Federated Shapley values. Wang et al. (Wang, 2019) discusses interpreting models in the context of Vertical Federated Learning (VFL) (Liu et al., 2022b) where different parties possess different slices of the feature space. Traditional model interpretation methods like Shapley values can reveal sensitive data across parties, making it unsuitable for VFL. To address this, a variant called SHAP Federated is proposed for VFL, particularly for dual-party scenarios involving a host and guest. The host and guest collaboratively develop a machine learning model, with the host owning the label data and part of the feature space, and the guest owning another part. The algorithm involves setting values in the instance $x$ to their original or reference values based on whether a feature is hosted or federated and encrypting IDs when necessary to maintain privacy. Then, predictions are made for each combination of features, and feature importance is calculated from the aggregated prediction results using Shapley values. Features that cannot handle missing values are set to either NA or the median (Lundberg and Lee, 2017).
@@ 5.3. Defences with Privacy-preserving ML models
Liu et al. (Liu et al., 2024b) develops a model-level defense by employing Differentially-Private Stochastic Gradient Descent (DP-SGD) (Bu et al., 2023), to build inherently private models. The process involves automatic configuration of gradient clipping and the selection of ‘MixOpt’ as the clipping model, uniformly applied across all model layers. While DP-SGD can reduce the effectiveness of membership inference attacks, it also significantly decreases classification accuracy, even with a large epsilon $\epsilon$. Findings indicate that attribution maps become less informative than even methods not considering model parameters (Hooker et al., 2019). This underscores the challenge of balancing between defense capability and performance utility, as effective defense mechanisms like DP-SGD can significantly impact model accuracy and the quality of explanations provided.
@@ 5.4. Defences with Perturbations
Jia et al. (Jia et al., 2019) introduces a defence technique called MemGuard, differing from other strategies that modify the training process. MemGuard cleverly injects perturbations into the confidence scores produced by the model for each input, transforming these altered scores into adversarial examples aimed at misleading attack models. However, the primary limitation of MemGuard is its focus on distorting the model’s output by adding noise, which does not protect the attribution maps, thus failing to completely deter the attacks (Liu et al., 2024b).
@@ 5.5. Defences with Anonymisation
Montenegro et al. (Montenegro et al., 2021) develops a privacy-preserving network with multi-class identity recognition designed for case-based explanations. The network seeks to preserve privacy by promoting a uniform distribution across identities, making identity recognition akin to random guessing. The PPRL-VGAN model (Chen et al., 2018a) (see 19(a)), which intentionally collapses to the replacement identity and task-related class, is replaced with a WGAN-GP framework that uses a Wasserstein loss with a gradient penalty to stabilise the discriminator (see 19(b)). This change, alongside using interpretability saliency maps for reconstruction of relevant task-related features, aims to retain the explanatory value in the privatised images (Montavon et al., 2017). Montenegro et al. (Montenegro et al., 2021) also introduces another privacy-preserving network that utilises a Siamese identity recognition framework to enhance privacy in domains with scarce images per subject. They employ a contrastive loss function for training, defined as $\text{ContrastiveLoss}=\frac{1}{2}\times Y\times ED^{2}+\frac{1}{2}\times(1-Y)\times[\max(0,m-ED)]^{2}$, where $Y$ is the label indicating if the image pair is of the same identity, $ED$ is the Euclidean distance between embeddings, and $m$ is a margin. The Siamese network ensures the privatised image is distinct in identity from the original and others in the dataset.
Privatised Counterfactual Samples. Montenegro et al. (Montenegro et al., 2021) also generates counterfactual explanations from the privatised samples. As shown in 19(c), a counterfactual generation module, in the form of a decoder, is added to the above privacy-preserving network to map an image’s latent representation to its counterfactual. This decoder is designed to make minimal alterations to the privatised factual explanations to change their predicted class, thereby minimising the pixel-wise distance between the factual and counterfactual explanations while altering the image’s task-related prediction. Saliency masks and explanatory features are used to guide changes to image regions that are relevant to the explanation. The loss function for the counterfactual decoder training is represented as $L_{C}=E_{I,M\sim p_{data}}[\lambda_{x}[F(I)\times(1-M)-C(I)\times(1-M)]^{2}+\lambda_{D}Exp(D_{exp}(I)\times\log(1-D_{exp}(C(I)))]$, where $F(I)$ and $C(I)$ denote the privatized factual and counterfactual explanations, respectively, and $\lambda_{x}$ and $\lambda_{D}$ are weights controlling the importance of each term in the loss function.
@@ 6.2. Published Datasets
Category Dataset #Items Disk Size Downstream Explanations Experimented in URL Image MNIST 70K 11MB Counterfactuals, Gradient (Huang et al., 2023; Yang et al., 2022; Zhao et al., 2021b; Milli et al., 2019) www.kaggle.com/datasets/hojjatk/mnist-dataset CIFAR 60K 163MB Gradient (Miura et al., 2021; Shokri et al., 2021; Milli et al., 2019; Liu et al., 2024b) www.cs.toronto.edu/~kriz/cifar.html SVHN 600K 400MB+ Gradient (Miura et al., 2021) ufldl.stanford.edu/housenumbers/ Food101 100K+ 10GB Case-based (Gaudio et al., 2023) www.kaggle.com/datasets/dansbecker/food-101 Flowers102 8K+ 300MB+ Case-based (Gaudio et al., 2023) www.robots.ox.ac.uk/~vgg/data/flowers/102/ Cervical 8K+ 46GB+ Case-based, Interpretable Models (Gaudio et al., 2023) www.kaggle.com/competitions/intel-mobileodt-cervical-cancer-screening CheXpert 220K+ GBs Case-based, Interpretable Models (Gaudio et al., 2023) stanfordmlgroup.github.io/competitions/chexpert/ Facial Expression 12K+ 63MB Black-box (Patel et al., 2022) www.kaggle.com/datasets/msambare/fer2013 Celeb 200K GBs Gradient (Zhao et al., 2021b) mmlab.ie.cuhk.edu.hk/projects/CelebA.html Tabular Adult 48K+ 10MB Counterfactuals, Shapley, Gradient, Perturbation 10+ ((Huang et al., 2023; Ferry et al., 2023a; Pentyala et al., 2023) etc.) archive.ics.uci.edu/ml/datasets/adult COMPAS 7K+ 25MB Counterfactuals, Shapley, Gradient, Perturbation (Ferry et al., 2023a; Duddu and Boutet, 2022) www.kaggle.com/datasets/danofer/compass FICO 10K+ $\leq$ 1MB Counterfactuals, Shapley (Huang et al., 2023; Wang et al., 2022; Pentyala et al., 2023; Pawelczyk et al., 2023) community.fico.com/s/explainable-machine-learning-challenge Boston Housing 500+ $\leq$ 1MB Counterfactuals, Shapley (Wang et al., 2022) www.kaggle.com/code/prasadperera/the-boston-housing-dataset German Credit 1K $\leq$ 1MB Counterfactuals, Shapley, Gradient, Perturbation (Vo et al., 2023; Goethals et al., 2023; Yang et al., 2022; Duddu and Boutet, 2022) archive.ics.uci.edu/dataset/144/statlog+german+credit+data Student Admission 500 $\leq$ 1MB Counterfactuals, Shapley (Vo et al., 2023) www.kaggle.com/datasets/mohansacharya/graduate-admissions Student Performance 10K $\leq$ 1MB Counterfactuals, Shapley (Vo et al., 2023) www.kaggle.com/datasets/nikhil7280/student-performance-multiple-linear-regression GMSC 150K+ 15MB Counterfactuals, Shapley (Wang et al., 2022; Naretto et al., 2022) www.kaggle.com/c/GiveMeSomeCredit/data Diabetes 100K+ 20MB Counterfactuals, Shapley (Pawelczyk et al., 2023; Luo et al., 2022; Yang et al., 2022; Watson et al., 2022; Shokri et al., 2021) archive.ics.uci.edu/dataset/296/diabetes+130-us+hospitals+for+years+1999-2008 Breast Cancer 569 $<1MB$ Interpretable models, Counterfactuals (Mochaourab et al., 2021) archive.ics.uci.edu/ml/datasets/breast+cancer Graph Cora 2K+ 4.5MB Feature-based (Olatunji et al., 2023) relational.fit.cvut.cz/dataset/CORA Bitcoin 30K $\leq$ 1MB Feature-based (Olatunji et al., 2023) snap.stanford.edu/data/soc-sign-bitcoin-alpha.html CIC-IDS2017 2.8M+ 500MB Counterfactuals (Kuppa and Le-Khac, 2021) www.unb.ca/cic/datasets/ids-2017.html Text IMDB Review 50K 66MB Black-box (Patel et al., 2022) ai.stanford.edu/~amaas/data/sentiment/
@@
The Bitcoin dataset (Kumar et al., 2016) is a network representation of trading accounts within the Bitcoin ecosystem (Liu et al., 2019). In this dataset, each trading account is depicted as a node, and there are weighted edges connecting pairs of accounts, symbolizing the level of trust between them. The weights range from +10, indicating complete trust, to -10, signifying complete distrust. Each node is labeled to denote its trustworthiness status. The feature vector associated with each node is derived from ratings provided by other users, including metrics such as average positive or negative ratings.
@@ 6.3. Evaluation Metrics
Category Evaluation Metrics Formula/Description Usage Explanation Utility Counterfactual validity (Goethals et al., 2023) Pureness=# value combinations with desired outcome# value $\text{Pureness}=\frac{\#\text{ value combinations with desired outcome}}{\#\text{ value combinations}}$ Assess the range of attribute values within k-anonymous counterfactual instances. Consider all attributes, including those beyond quasi-identifiers Classification metric (Goethals et al., 2023) $CM=\frac{\sum_{i=1}^{N}\text{penalty}(tuple_{i})}{N}$ Assess equivalence classes within anonymized datasets, focusing on class label uniformity. Faithfulness (RDT-Fidelity) (Olatunji et al., 2023; Funke et al., 2022) $\mathcal{F}(\mathcal{E}_{X})=\mathbb{E}_{Y_{\mathcal{E}_{X}}|Z\sim\mathcal{N}}\left[1_{f(X)=f(Y_{\mathcal{E}_{X}})}\right]$ Reflect how often the model’s predictions are unchanged despite perturbations to the input, which would suggest that the explanation is effectively capturing the reasoning behind the model’s predictions. Sparsity (Olatunji et al., 2023; Funke et al., 2022) $H(p)=-\sum_{f\in M}p(f)\log p(f)$ A complete and faithful explanation to the model should inherently be sparse, focusing only on a select subset of features that are most predictive of the model’s decision. Information Loss Normalised Certainty Penalty (NCP) (Goethals et al., 2023) $\text{NCP}(G)=\sum_{i=1}^{d}w_{i}\cdot\text{NCP}_{A_{i}}(G)$ Higher NCP values indicate a greater degree of generalization and more information loss. This metric helps in assessing the balance between data privacy and utility. Discernibility (Goethals et al., 2023) $C_{DM}(g,k)=\sum_{VE\,s.t.\,|E|\geq k}|E|^{2}+\sum_{VE\,s.t.\,|E|<k}|D||E|$ Measure the penalties on tuples in a dataset after k-anonymization, reflecting how indistinguishable they are post-anonymization Approximation Loss (Goethals et al., 2023) $\mathcal{E}(\hat{\phi},\mathcal{Z},f(X))\triangleq\mathbb{E}[\mathcal{L}(\hat{\phi},\mathcal{Z},f(X))-\mathcal{L}(\phi^{*},\mathcal{Z},f(X))].$ Measure the error caused by randomness added when minimizing the privacy loss as the expected deviation of the randomized explanation from the best local approximation Explanation Intersection (Olatunji et al., 2023; Funke et al., 2022) The percentage of bits in the original explanation that is retained in the privatised explanation after using differential privacy The higher the better but due to privacy-utility trade-off, this metric should not be 100%. Privacy Degree $k$-anonymity (Goethals et al., 2023) A person’s information is indistinguishable from at least k-1 other individuals. Refers to the number of individuals in the training dataset to whom a given explanation could potentially be linked (Goethals et al., 2023). Information Leakage (Patel et al., 2022) $Pr_{i=1..k}\hat{\phi}(\mathbf{z_{i}},X,f_{D}(X))\leq e^{\hat{\varepsilon}}\cdot Pr[\hat{\phi}(\mathbf{z_{i}},X,f^{\prime}_{D}(X)):\forall i]+\hat{\delta}$ If an adversary can access model explanations, they would not gain any additional information that could help in inferring something about the training data beyond what could be learned from the model predictions alone Privacy Budget The total privacy budget for all queries is fixed at ($\varepsilon,\delta$). The explanation algorithm must not exceed the overall budget across all queries. Stricter requirement ($\varepsilon_{min},\delta_{min}$) is set for each individual query. Attack Success Precision/Recall/F1 (Duddu and Boutet, 2022) $Prec=\frac{TP}{TP+FP}$, $Rec=\frac{TP}{TP+FN}$, $F1=2\times\frac{\text{precision}\times\text{recall}}{\text{precision}+\text{recall}}$ Evaluate an attack’s effectiveness in correctly and completely identifying the properties it is designed to infer. Balanced Accuracy (Liu et al., 2024b; Pawelczyk et al., 2023; Huang et al., 2023) $BA=\frac{TPR+TNR}{2}$ Measures the accuracy of attack (e.g. membership prediction in membership inference attacks), on a balanced dataset of members and non-members. ROC/AUC (Huang et al., 2023; Pawelczyk et al., 2023; Liu et al., 2024b; Ferry et al., 2023a; Olatunji et al., 2023) The ROC curve plots the true positive rate against the false positive rate at various threshold settings. An AUC near 1 indicates a highly successful privacy attack, while an AUC close to 0.5 suggests no better performance than random guessing. TPR at Low FPR (Liu et al., 2024b; Huang et al., 2023; Pawelczyk et al., 2023) Report TPR at a fixed FPR (e.g., 0.1%). If an attack can pinpoint even a minuscule fraction of the training dataset with high precision, then the attack ought to be deemed effective. Mean Absolute Error (MAE) (Luo et al., 2022) $\ell_{1}(\hat{x},x)=\frac{1}{mn}\sum_{j=1}^{m}\sum_{i=1}^{n}|\hat{x}_{i}^{j}-x_{i}^{j}|,$ Gives an overview of how accurately an attack can reconstruct private inputs by averaging the absolute differences across all samples and features. Success Rate (SR) (Luo et al., 2022) $SR=\frac{|\hat{X}_{val}\neq\perp|}{mn}$ The ratio of successfully reconstructed features to the total number of features across all samples Model Agreement (Wang et al., 2022) $\text{Agreement}=\frac{1}{n}\sum_{i=1}^{n}1_{f_{\theta}(x_{i})=h_{\phi}(x_{i})}.$ A higher agreement indicates that the substitute model is more similar to the original model. When comparing two model extraction methods with the same agreement, the one with the lower standard deviation is preferred. Average Uncertainty Reduction (Ferry et al., 2023a) $Dist(\mathcal{D}^{M},\mathcal{D}^{Orig})=\frac{1}{n\cdot d}\sum_{i=1}^{n}\sum_{k=1}^{d}\frac{H(\mathcal{D}^{M}_{i,k})}{H(\mathcal{D}_{i,k})}$ The degree to which a data reconstruction attack is accurate, measured by the reduction in uncertainty across all features of all samples in the dataset

### v2

@@ 5.2. Defences with Privacy-Preserving SHAP
Several studies have focused on preserve the privacy of users from explanation using Shapley values, including quatization, dimension reduction, multi-party computation, federated learning, and differential privacy (see Fig. 12).
@@
Federated Shapley values. Wang et al. (Wang, 2019) discusses interpreting models in the context of Vertical Federated Learning (VFL) (Liu et al., 2024a; Liu et al., 2023; Liu et al., 2022b, 2024b) where different parties possess different slices of the feature space. Traditional model interpretation methods like Shapley values can reveal sensitive data across parties, making it unsuitable for VFL. To address this, a variant called SHAP Federated is proposed for VFL, particularly for dual-party scenarios involving a host and guest. The host and guest collaboratively develop a machine learning model, with the host owning the label data and part of the feature space, and the guest owning another part. The algorithm involves setting values in the instance $x$ to their original or reference values based on whether a feature is hosted or federated and encrypting IDs when necessary to maintain privacy. Then, predictions are made for each combination of features, and feature importance is calculated from the aggregated prediction results using Shapley values. Features that cannot handle missing values are set to either NA or the median (Lundberg and Lee, 2017).
@@ 5.3. Defences with Privacy-preserving ML models
Liu et al. (Liu et al., 2024d) develops a model-level defense by employing Differentially-Private Stochastic Gradient Descent (DP-SGD) (Bu et al., 2023), to build inherently private models. The process involves automatic configuration of gradient clipping and the selection of ‘MixOpt’ as the clipping model, uniformly applied across all model layers. While DP-SGD can reduce the effectiveness of membership inference attacks, it also significantly decreases classification accuracy, even with a large epsilon $\epsilon$. Findings indicate that attribution maps become less informative than even methods not considering model parameters (Hooker et al., 2019). This underscores the challenge of balancing between defense capability and performance utility, as effective defense mechanisms like DP-SGD can significantly impact model accuracy and the quality of explanations provided.
@@ 5.4. Defences with Perturbations
Jia et al. (Jia et al., 2019b) introduces a defence technique called MemGuard, differing from other strategies that modify the training process. MemGuard cleverly injects perturbations into the confidence scores produced by the model for each input, transforming these altered scores into adversarial examples aimed at misleading attack models. However, the primary limitation of MemGuard is its focus on distorting the model’s output by adding noise, which does not protect the attribution maps, thus failing to completely deter the attacks (Liu et al., 2024d).
@@ 5.5. Defences with Anonymisation
Montenegro et al. (Montenegro et al., 2021) develops a privacy-preserving network with multi-class identity recognition designed for case-based explanations. The network seeks to preserve privacy by promoting a uniform distribution across identities, making identity recognition akin to random guessing. The PPRL-VGAN model (Chen et al., 2018a) (see 16(a)), which intentionally collapses to the replacement identity and task-related class, is replaced with a WGAN-GP framework that uses a Wasserstein loss with a gradient penalty to stabilise the discriminator (see 16(b)). This change, alongside using interpretability saliency maps for reconstruction of relevant task-related features, aims to retain the explanatory value in the privatised images (Montavon et al., 2017). Montenegro et al. (Montenegro et al., 2021) also introduces another privacy-preserving network that utilises a Siamese identity recognition framework to enhance privacy in domains with scarce images per subject. They employ a contrastive loss function for training, defined as $\text{ContrastiveLoss}=\frac{1}{2}\times Y\times ED^{2}+\frac{1}{2}\times(1-Y)\times[\max(0,m-ED)]^{2}$, where $Y$ is the label indicating if the image pair is of the same identity, $ED$ is the Euclidean distance between embeddings, and $m$ is a margin. The Siamese network ensures the privatised image is distinct in identity from the original and others in the dataset.
Privatised Counterfactual Samples. Montenegro et al. (Montenegro et al., 2021) also generates counterfactual explanations from the privatised samples. As shown in 16(c), a counterfactual generation module, in the form of a decoder, is added to the above privacy-preserving network to map an image’s latent representation to its counterfactual. This decoder is designed to make minimal alterations to the privatised factual explanations to change their predicted class, thereby minimising the pixel-wise distance between the factual and counterfactual explanations while altering the image’s task-related prediction. Saliency masks and explanatory features are used to guide changes to image regions that are relevant to the explanation. The loss function for the counterfactual decoder training is represented as $L_{C}=E_{I,M\sim p_{data}}[\lambda_{x}[F(I)\times(1-M)-C(I)\times(1-M)]^{2}+\lambda_{D}Exp(D_{exp}(I)\times\log(1-D_{exp}(C(I)))]$, where $F(I)$ and $C(I)$ denote the privatized factual and counterfactual explanations, respectively, and $\lambda_{x}$ and $\lambda_{D}$ are weights controlling the importance of each term in the loss function.
@@ 6.2. Published Datasets
Category Dataset #Items Disk Size Downstream Explanations Experimented in URL Image MNIST 70K 11MB Counterfactuals, Gradient (Huang et al., 2023; Yang et al., 2022; Zhao et al., 2021b; Milli et al., 2019) www.kaggle.com/datasets/hojjatk/mnist-dataset CIFAR 60K 163MB Gradient (Miura et al., 2021; Shokri et al., 2021; Milli et al., 2019; Liu et al., 2024d) www.cs.toronto.edu/~kriz/cifar.html SVHN 600K 400MB+ Gradient (Miura et al., 2021) ufldl.stanford.edu/housenumbers/ Food101 100K+ 10GB Case-based (Gaudio et al., 2023) www.kaggle.com/datasets/dansbecker/food-101 Flowers102 8K+ 300MB+ Case-based (Gaudio et al., 2023) www.robots.ox.ac.uk/~vgg/data/flowers/102/ Cervical 8K+ 46GB+ Case-based, Interpretable Models (Gaudio et al., 2023) www.kaggle.com/competitions/intel-mobileodt-cervical-cancer-screening CheXpert 220K+ GBs Case-based, Interpretable Models (Gaudio et al., 2023) stanfordmlgroup.github.io/competitions/chexpert/ Facial Expression 12K+ 63MB Black-box (Patel et al., 2022) www.kaggle.com/datasets/msambare/fer2013 Celeb 200K GBs Gradient (Zhao et al., 2021b) mmlab.ie.cuhk.edu.hk/projects/CelebA.html Tabular Adult 48K+ 10MB Counterfactuals, Shapley, Gradient, Perturbation 10+ ((Huang et al., 2023; Ferry et al., 2023a; Pentyala et al., 2023) etc.) archive.ics.uci.edu/ml/datasets/adult COMPAS 7K+ 25MB Counterfactuals, Shapley, Gradient, Perturbation (Ferry et al., 2023a; Duddu and Boutet, 2022) www.kaggle.com/datasets/danofer/compass FICO 10K+ $\leq$ 1MB Counterfactuals, Shapley (Huang et al., 2023; Wang et al., 2022; Pentyala et al., 2023; Pawelczyk et al., 2023) community.fico.com/s/explainable-machine-learning-challenge Boston Housing 500+ $\leq$ 1MB Counterfactuals, Shapley (Wang et al., 2022) www.kaggle.com/code/prasadperera/the-boston-housing-dataset German Credit 1K $\leq$ 1MB Counterfactuals, Shapley, Gradient, Perturbation (Vo et al., 2023; Goethals et al., 2023; Yang et al., 2022; Duddu and Boutet, 2022) archive.ics.uci.edu/dataset/144/statlog+german+credit+data Student Admission 500 $\leq$ 1MB Counterfactuals, Shapley (Vo et al., 2023) www.kaggle.com/datasets/mohansacharya/graduate-admissions Student Performance 10K $\leq$ 1MB Counterfactuals, Shapley (Vo et al., 2023) www.kaggle.com/datasets/nikhil7280/student-performance-multiple-linear-regression GMSC 150K+ 15MB Counterfactuals, Shapley (Wang et al., 2022; Naretto et al., 2022) www.kaggle.com/c/GiveMeSomeCredit/data Diabetes 100K+ 20MB Counterfactuals, Shapley (Pawelczyk et al., 2023; Luo et al., 2022; Yang et al., 2022; Watson et al., 2022; Shokri et al., 2021) archive.ics.uci.edu/dataset/296/diabetes+130-us+hospitals+for+years+1999-2008 Breast Cancer 569 $<1MB$ Interpretable models, Counterfactuals (Mochaourab et al., 2021) archive.ics.uci.edu/ml/datasets/breast+cancer Graph Cora 2K+ 4.5MB Feature-based (Olatunji et al., 2023) relational.fit.cvut.cz/dataset/CORA Bitcoin 30K $\leq$ 1MB Feature-based (Olatunji et al., 2023) snap.stanford.edu/data/soc-sign-bitcoin-alpha.html CIC-IDS2017 2.8M+ 500MB Counterfactuals (Kuppa and Le-Khac, 2021) www.unb.ca/cic/datasets/ids-2017.html Text IMDB Review 50K 66MB Black-box (Patel et al., 2022) ai.stanford.edu/~amaas/data/sentiment/
@@
The Bitcoin dataset (Kumar et al., 2016) is a network representation of trading accounts within the Bitcoin ecosystem. In this dataset, each trading account is depicted as a node, and there are weighted edges connecting pairs of accounts, symbolizing the level of trust between them. The weights range from +10, indicating complete trust, to -10, signifying complete distrust. Each node is labeled to denote its trustworthiness status. The feature vector associated with each node is derived from ratings provided by other users, including metrics such as average positive or negative ratings.
@@ 6.3. Evaluation Metrics
Category Evaluation Metrics Formula/Description Usage Explanation Utility Counterfactual validity (Goethals et al., 2023) Pureness=# value combinations with desired outcome# value $\text{Pureness}=\frac{\text{\# value combinations with desired outcome}}{\text{\# value combinations}}$ Assess the range of attribute values within k-anonymous counterfactual instances. Consider all attributes, including those beyond quasi-identifiers Classification metric (Goethals et al., 2023) $CM=\frac{\sum_{i=1}^{N}\text{penalty}(tuple_{i})}{N}$ Assess equivalence classes within anonymized datasets, focusing on class label uniformity. Faithfulness (RDT-Fidelity) (Olatunji et al., 2023; Funke et al., 2022) $\mathcal{F}(\mathcal{E}_{X})=\mathbb{E}_{Y_{\mathcal{E}_{X}}|Z\sim\mathcal{N}}\left[1_{f(X)=f(Y_{\mathcal{E}_{X}})}\right]$ Reflect how often the model’s predictions are unchanged despite perturbations to the input, which would suggest that the explanation is effectively capturing the reasoning behind the model’s predictions. Sparsity (Olatunji et al., 2023; Funke et al., 2022) $H(p)=-\sum_{f\in M}p(f)\log p(f)$ A complete and faithful explanation to the model should inherently be sparse, focusing only on a select subset of features that are most predictive of the model’s decision. Information Loss Normalised Certainty Penalty (NCP) (Goethals et al., 2023) $\text{NCP}(G)=\sum_{i=1}^{d}w_{i}\cdot\text{NCP}_{A_{i}}(G)$ Higher NCP values indicate a greater degree of generalization and more information loss. This metric helps in assessing the balance between data privacy and utility. Discernibility (Goethals et al., 2023) $C_{DM}(g,k)=\sum_{VE\,s.t.\,|E|\geq k}|E|^{2}+\sum_{VE\,s.t.\,|E|<k}|D||E|$ Measure the penalties on tuples in a dataset after k-anonymization, reflecting how indistinguishable they are post-anonymization Approximation Loss (Goethals et al., 2023) $\mathcal{E}(\hat{\phi},\mathcal{Z},f(X))\triangleq\mathbb{E}[\mathcal{L}(\hat{\phi},\mathcal{Z},f(X))-\mathcal{L}(\phi^{*},\mathcal{Z},f(X))].$ Measure the error caused by randomness added when minimizing the privacy loss as the expected deviation of the randomized explanation from the best local approximation Explanation Intersection (Olatunji et al., 2023; Funke et al., 2022) The percentage of bits in the original explanation that is retained in the privatised explanation after using differential privacy The higher the better but due to privacy-utility trade-off, this metric should not be 100%. Privacy Degree $k$-anonymity (Goethals et al., 2023) A person’s information is indistinguishable from at least k-1 other individuals. Refers to the number of individuals in the training dataset to whom a given explanation could potentially be linked (Goethals et al., 2023). Information Leakage (Patel et al., 2022) $Pr_{i=1..k}\hat{\phi}(\mathbf{z_{i}},X,f_{D}(X))\leq e^{\hat{\varepsilon}}\cdot Pr[\hat{\phi}(\mathbf{z_{i}},X,f^{\prime}_{D}(X)):\forall i]+\hat{\delta}$ If an adversary can access model explanations, they would not gain any additional information that could help in inferring something about the training data beyond what could be learned from the model predictions alone Privacy Budget The total privacy budget for all queries is fixed at ($\varepsilon,\delta$). The explanation algorithm must not exceed the overall budget across all queries. Stricter requirement ($\varepsilon_{min},\delta_{min}$) is set for each individual query. Attack Success Precision/Recall/F1 (Duddu and Boutet, 2022) $Prec=\frac{TP}{TP+FP}$, $Rec=\frac{TP}{TP+FN}$, $F1=2\times\frac{\text{precision}\times\text{recall}}{\text{precision}+\text{recall}}$ Evaluate an attack’s effectiveness in correctly and completely identifying the properties it is designed to infer. Balanced Accuracy (Liu et al., 2024d; Pawelczyk et al., 2023; Huang et al., 2023) $BA=\frac{TPR+TNR}{2}$ Measures the accuracy of attack (e.g. membership prediction in membership inference attacks), on a balanced dataset of members and non-members. ROC/AUC (Huang et al., 2023; Pawelczyk et al., 2023; Liu et al., 2024d; Ferry et al., 2023a; Olatunji et al., 2023) The ROC curve plots the true positive rate against the false positive rate at various threshold settings. An AUC near 1 indicates a highly successful privacy attack, while an AUC close to 0.5 suggests no better performance than random guessing. TPR at Low FPR (Liu et al., 2024d; Huang et al., 2023; Pawelczyk et al., 2023) Report TPR at a fixed FPR (e.g., 0.1%). If an attack can pinpoint even a minuscule fraction of the training dataset with high precision, then the attack ought to be deemed effective. Mean Absolute Error (MAE) (Luo et al., 2022) $\ell_{1}(\hat{x},x)=\frac{1}{mn}\sum_{j=1}^{m}\sum_{i=1}^{n}|\hat{x}_{i}^{j}-x_{i}^{j}|,$ Gives an overview of how accurately an attack can reconstruct private inputs by averaging the absolute differences across all samples and features. Success Rate (SR) (Luo et al., 2022) $SR=\frac{|\hat{X}_{val}\neq\perp|}{mn}$ The ratio of successfully reconstructed features to the total number of features across all samples Model Agreement (Wang et al., 2022) $\text{Agreement}=\frac{1}{n}\sum_{i=1}^{n}1_{f_{\theta}(x_{i})=h_{\phi}(x_{i})}.$ A higher agreement indicates that the substitute model is more similar to the original model. When comparing two model extraction methods with the same agreement, the one with the lower standard deviation is preferred. Average Uncertainty Reduction (Ferry et al., 2023a) $Dist(\mathcal{D}^{M},\mathcal{D}^{Orig})=\frac{1}{n\cdot d}\sum_{i=1}^{n}\sum_{k=1}^{d}\frac{H(\mathcal{D}^{M}_{i,k})}{H(\mathcal{D}_{i,k})}$ The degree to which a data reconstruction attack is accurate, measured by the reduction in uncertainty across all features of all samples in the dataset